# Whispers in the Machine: Confidentiality in LLM-integrated Systems

### v3

@@ Abstract
Large Language Models (LLMs) are increasingly augmented with external tools and commercial services into LLM-integrated systems. While these interfaces can significantly enhance the capabilities of the models, they also introduce a new attack surface. Manipulated integrations, for example, can exploit the model and compromise sensitive data accessed through other interfaces. While previous work primarily focused on attacks targeting a model’s alignment or the leakage of training data, the security of data that is only available during inference has escaped scrutiny so far. In this work, we demonstrate the vulnerabilities associated with external components and introduce a systematic approach to evaluate confidentiality risks in LLM-integrated systems. We identify two specific attack scenarios unique to these systems and formalize these into a tool-robustness framework designed to measure a model’s ability to protect sensitive information. Our findings show that all examined models are highly vulnerable to confidentiality attacks, with the risk increasing significantly when models are used together with external tools.
@@
Large language models, Confidentiality attacks, Machine learning, Prompt injection attacks, LLM-integrated systems
@@ 1 Introduction
Figure 1: Confidentiality in LLM-integrated systems. We consider systems where the user is able to interact with services and functionalities via an internal large language model which processes natural language instructions. The model is initialized with a set of instructions and—to extend its capabilities—has access to external services and tools through clearly defined integrations. An attacker can compromise the system exploiting vulnerabilities in the integrated language model by either acting as the user or by insert malicious data into the used services. Prompt-based attacks inside the users input or malicious instruction inside the accessed services and tools can lead to extraction of sensitive information stored in the external services and tools.
The capabilities of Large Language Models (LLMs) can be significantly enhanced through integrations with external systems. These integrations range from incorporating LLMs with organizational knowledge bases [1] to enhancing business tools like email [2], cloud, and calendar [3] services. Notably, Google and Apple have recently announced to include LLMs into their operating systems as digital assistants representing an instance of a LLM-integrated system [4, 5].
@@
While a broad range of integrations can greatly expand the capabilities of an LLM-integrated system, it also increases its attack surface. When language models are embedded into real-world applications, they may access sensitive information, making data security dependent on how the model and its orchestration manage and protect this data. Without robust safeguards, a language model may inadvertently expose internal or personal information, introducing potential security risks.
This is not only a theoretical risk but illustrated by recent examples. For instance, Microsoft’s Copilot [6]—an AI-assistant based on GPT-4 build for the current versions of Windows—was exploited in a recent attack where a malicious email initiated a prompt-based exploit to extract a user’s personal information [7]. This shows how vulnerabilities in an integration can potentially compromise the entire system. Indeed, according to the OWASP Foundation, insecure plugin and integration design is one of the ten most significant vulnerabilities in the context of LLMs. Moreover, in the Large Language Model Capture-the-Flag (LLM CTF) competition at SaTML 2024, where participants aim to protect their own secrets while stealing the secrets of other teams, all proposed defenses were found to be vulnerable against at least one attack [8]. This reinforces the motivation to take a closer look at the robustness within LLM integrations.
@@
So far, the research community has mainly concentrated on attacks aimed at a model’s alignment [9, 10, 11, 12] or on the leakage of training data [13], while the security of data that is available only during inference has escaped scrutiny. In this work, we address this gap by systematically studying confidentiality in LLM-integrated systems. Figure 1 illustrates a typical setup of such a system in which a LLM has access to various integrations, such as email, calendars, notes, and cloud services. While the model is authorized to access data from these sources for internal use, the potential disclosure of sensitive information between integrations poses a security risk. Our primary objectives are to systematize the attack surface associated with integrating LLMs and to explore their ability to maintain the confidentiality of sensitive information. This is challenging as sensitive data in these systems is complex and often ambiguous. While certain types of information, like passwords and encryption keys, are almost universally considered sensitive, other data—such as files in formats like PDFs—may contain personal details (e.g., banking or health information). This variability makes it difficult to apply strict rules, and may require users to establish their own data protection policies.
@@
To account for this, we first establish a “secret-key” game that captures the (in-)ability of the model to understand confidentiality. In this game, we use a secret key as a proxy for confidential data, which could be accessed through integrations. The goal for the attacker is to prompt the model in such a way that the secret key can be extracted from the model’s output. Under this definition, we can accurately measure the robustness of a model and compare 14 different attack strategies and 5 defense mechanisms including the recently proposed PromptGuard from Meta [14]. Building upon this, we then expand the game into a “tool-robustness” framework, where the language model operates within a system equipped with external tool integrations. For this, we define two primary threat scenarios: one in which the adversary acts as a malicious integration and another where a user account is compromised. These threats are explored through 20 scenarios that combine different tools, allowing us to evaluate the overall robustness of language models in systems integrated with external tools.
@@
Previous research has demonstrated the general vulnerability of language models to prompt-based attacks and our findings reveal an even higher susceptibility in systems where these models are integrated with external tools. Based on our tool-robustness framework, we show that all considered LLMs have an increased vulnerability to data extraction when integrated into systems with external tools. The overall success rate of confidential data extraction is increased by up to 73 % for single-tool usage and by up to 49 % for scenarios with multi-tool usage.
@@
In summary, we find that embedding LLMs into real-world systems poses a fundamental risk to user privacy and security. Refinement through fine-tuning can improve model performance, the inherent risks remain due to the fundamental nature of these models. This highlights the need for continued research and carefully designed systems to address these challenges effectively.
@@
  • Confidentiality in LLM-integrated systems. We demonstrate the risks of integrating LLMs with external services and tools. We consider two primary threat scenarios in which the adversary acts either as a malicious integration or compromised user.
  • Tool-robustness framework. We propose a formal framework to analyze the confidentiality of LLM-integrated systems. The framework consists of two main components: First, the “secret-key game”, which initializes the LLM with a secret key and a set of safeguards. Second, we extend the game into a “tool-robustness” framework, where the language model operates within a system equipped with external tool integrations. In this framework, we define 20 unique scenarios of tool interactions to measure the robustness of LLMs in integrated systems with external tools.
  • LLM Evaluation. Within our framework, we extensively evaluate 6 state-of-the-art open-source LLMs, ranging in size from 1 billion parameters for on-device usage to 72 billion for extensive capabilities. We test our scenarios on a dataset of 2,000 system prompts containing safeguards designed to guide model behavior. We also examine how attacks and defenses from our “secret-key” game impact model alignment in tool-augmented contexts.
All code, the generated and used datasets, and instructions on how to reproduce our results are published at: https://github.com/LostOxygen/llm-confidentiality.
@@ 2 Technical Background
To enable adaptability across various tasks, the models are often refined during training to follow so-called system prompts. The basic idea is to prefix an input prompt with an instruction that adapts a model’s alignment without further training. Therefore, the model is fine-tuned on specific role-based data. This contains specific tokens to distinguish between different roles and allows the model to differ between system prompt instructions and user-supplied inputs. System prompts are widely used to align state-of-the-art LLMs with specific use cases. For example, the following template is used to fine-tune the chat version of Meta’s Llama 3.1 model [16]:
@@
Prompt-based attacks. Although the predetermined system prompts may appear to be separate from the remaining data and user inputs within the input prompts,
the absence of system level separation between the high-priority instructions inside the system prompt and the user requested instructions inside the user input segments can lead to vulnerabilities exploiting these conditions. Figure 2 shows an example of a prompt injection attack: The system prompt $x^{sys}$ clearly indicates that the key 1337 is a secret and should not be part of the LLM response. However, the user input contains an alternative instruction-like text, which can be able to overwrite the system prompt if not detected.
@@
Figure 3: Data leakage attacks. Overview over the motivational example based on actual Google Mail and Google Drive integrations. The LLM is instructed by the user to summarize an email as in day-to-day life (Step ❶). The email however is manipulated by an attacker and contains malicious instructions to hijack the LLM (Step ❷). Without the user knowing from the malicious instructions, the LLM’s initial instructions get overwritten with the new instructions to instead access Google Drive, search for a file containing a secret key, and send it back to the attacker via email. As a result the LLM accesses a confidential file on Google Drive (Step ❸) and sends it back to the attacker via email (Step ❹) without the possibility for the user to intervene or even knowing about this.
@@
Prompt-based integration. This approach involves crafting specific prompts or instructions that guide the model to use a tool or service effectively. The language model interacts with external tools through these instructions, without altering the model’s underlying structure. It relies on the model’s ability to understand the task from the input prompt and communicate with the tool accordingly. This method is faster and more flexible since it doesn’t require modifying the model, and can be adapted for various tasks without retraining. A commonly used framework to integrate external services is ReAct [18]. ReAct synergizes reasoning and acting directions to guide the LLM into expanding its knowledge—if necessary for the task—via external services, e.g., performing database queries or assessing a calendar if asked to look for specific events. This results in fewer hallucinations, made up information by the model, since the model consults the appropriate service to receive more factual information to solve its task.
@@ 3 Confidentiality in LLM Integrations
We define LLM-integrated systems as systems where a language model serves as an interface between the user and the underlying system, facilitating user requests by integrating with external services and tools. An example of such a system is shown in Figure 1. These systems are primarily designed to assist users by providing access to information (such as from databases or knowledge bases) or by automating common tasks like scheduling meetings or sending emails.
@@
Motivational example. To illustrate the potential for data leakage, we begin with a concrete example. We consider a scenario in which the language model has access to both email client and a cloud service. First, the user asks the model to summarize their latest email as it would possible happen in day-to-day life. However, as shown in Figure 8, the email app is injected with a malicious email from an attacker containing instructions directing the model to access the cloud service and retrieve a specific file without the user’s knowledge. Without the user knowing of these instructions in the first place, the LLM’s initial instructions may be overwritten by the new ones. This leads the LLM to access the file from the cloud service and send it back to the attacker via email without the possibility for the user to intervene.
@@
The results are presented in Table I. We observe a high success rate, with 99 % of all attempts by the attacker to extract the secret key by hijacking the LLM. This initial experiment demonstrates the necessity of further safeguards to restrict the LLM’s behavior.
@@
We find that this approach is effective, reducing the secret leakage rate to 34 % for the 8b model and 21 % for the 70b model. This demonstrates that the model is capable of understanding the sensitivity of the data and can—in principle—be effectively guided by the use of appropriate safeguards.
@@
First, we consider translation attacks which translate malicious instructions into languages other than English and asking the language model to translate and execute them [21, 22]. This, allows the attacker to misalign the LLM from its original instructions. Second, we consider jailbreak attacks, which trick the LLM into doing normally restricted actions [23, 24, 25, 26]. An example email using the translation attack is shown in Figure 4. In this scenario, the success rate increases significantly again, reaching 97 % for the smaller and 94 % for the larger model.
@@
Scenario / LLM Llama 3.1 8b Llama 3.1 70b
Data extraction without safeguards 99 % 99 %
Data extraction with safeguards 34 % 21 %
Data extraction with bypass 97 % 94 %
@@ 4 Secret-key game
To this end, we start by establishing a clear definition of an attacker’s success. We formalize the ability of the model to “keep a secret” in a game executed between an LLM and an attacker, as depicted in Figure 5. In this game, we embed a secret string $s$ in the private system prompt $x^{sys}$ of the LLM together with instructions not to leak this secret. Figure 2 shows an example of such a prompt. In our game, the secret $s$ represents a general proxy for confidential information and the game serves as a foundational benchmark for assessing a model’s capability to keep information confidential—a model that cannot keep a secret in this game is unlikely to protect confidential information in the deployed systems.
@@
System prompts. While the LLM is always instructed to keep confidential data safe and the key secret, we accompany this system prompt with further instructions to enhance its functionality. To this end, create a dataset of functional system prompts. We begin with a set of manually crafted system prompts, then use a language model to generate variations of these prompts. The specific instruction prompt given to the language model is detailed in Appendix F. These enhanced system prompts are compiled into a dataset for use alongside the original system prompt. In total, we create 2000 different system prompts.
@@ 4.1 Attacks
Models / Scenarios Llama 3.1 Llama 3.1
@@
Payload Splitting 1 (+1) 0 (±$\pm 0$) 7.8 (+7.8) 9.6 (+9.6)
Obfuscation 0 (±$\pm 0$) 0 (±$\pm 0$) 0 (±$\pm 0$) 0 (±$\pm 0$)
@@
Typoglycemia 0 (±$\pm 0$) 0 (±$\pm 0$) 0 (±$\pm 0$) 0 (±$\pm 0$)
Adversarial Suffix 2 (+2) 0 (±$\pm 0$) 2.4 (+2.4) 2(+2)
Prefix Injection 0 (±$\pm 0$) 45 (+45) 1.2 (+1.2) 30.6 (+30.6)
@@
Context Ignoring 91 (+91) 81 (+81) 0 (±$\pm 0$) 36.6 (+36.6)
Context Termination 0 (+) 0 (±$\pm 0$) 39.6 (+39.6) 2 (+2)
@@
Few-Shot 0 (±$\pm 0$) 7 (+7) 0 (±$\pm 0$) 7.6 (+7.6)
Cognitive Hacking 0 (±$\pm 0$) 0 (±$\pm 0$) 0 (±$\pm 0$) 1.4 (+1.4)

### v4

# Whispers in the Machine: Confidentiality in Agentic Systems

@@ Abstract
The interaction between users and applications is increasingly shifted toward natural language by deploying Large Language Models (LLMs) as the core interface. The capabilities of these so-called agents become more capable the more tools and services they serve as an interface for, ultimately leading to agentic systems. Agentic systems use LLM-based agents as interfaces for most user interactions and various integrations with external tools and services. While these interfaces can significantly enhance the capabilities of the agentic system, they also introduce a new attack surface. Manipulated integrations, for example, can exploit the internal LLM and compromise sensitive data accessed through other interfaces. While previous work primarily focused on attacks targeting a model’s alignment or the leakage of training data, the security of data that is only available during inference has escaped scrutiny so far. In this work, we demonstrate how the integration of LLMs into systems with external tool integration poses a risk similar to established prompt-based attacks, able to compromise the confidentiality of the entire system. Introducing a systematic approach to evaluate these confidentiality risks, we identify two specific attack scenarios unique to these agentic systems and formalize these into a tool-robustness framework designed to measure a model’s ability to protect sensitive information. Our analysis reveals significant vulnerabilities across all tested models, highlighting an increased risk when models are combined with external tools.
@@
Large language models, Confidentiality attacks, Machine learning, Prompt injection attacks, Agents, Agentic systems
@@ 1 Introduction
Figure 1: Confidentiality in agentic systems. We consider agent-based systems in which a mostly autonomous agent serves as the primary interface for users interacting with services and functionalities using natural language instructions. The agent is powered by an large language model initialized with a set of instructions and—to extend its capabilities—access to external services and tools through clearly defined integrations. An attacker can compromise the system exploiting vulnerabilities in the language model by either acting as the user or by insert malicious data into the used services. Prompt-based attacks inside the users input or malicious instruction inside the accessed services and tools can lead to extraction of sensitive information stored in the external services and tools.
The interaction between users and applications can be significantly enhanced and simplified by deploying so-called agents based on Large Language Models (LLMs) as an interface. These agents do not only allow users to interact with an application via natural language. However, they can also be integrated with various tools and services to enhance their functionality. Integrations can span a wide range from connecting the LLM-based agents to organizational knowledge bases [1], productivity tools like email [2], cloud storage, or calendar services [3]. The more tools and services an agent is integrated with, the more capable it is, ultimately leading to extensive agentic systems where the agent serves as an interface for a whole ecosystem of services and tools. For this, Google and Apple have recently announced plans to embed LLMs directly into their operating systems as digital assistants, representing a real-world example of such an agentic system [4, 5].
@@
While a broad range of integrations can unlock powerful new capabilities, it also broadens the attack surface. When language models are integrated with real-world systems, they may gain access to sensitive data. The security of this data then depends on how well the model and its orchestration layers manage and protect it. Without robust safeguards, an LLM may inadvertently expose internal or personal information, introducing a significant security risk.
This risk is not merely theoretical but illustrated by recent examples. For example, Microsoft’s Copilot [6]—a GPT-4-powered assistant built into Windows—was exploited through a prompt-based attack triggered by a malicious email with the goal to extract a user’s personal information [7]. This demonstrates how LLM-driven integrations can serve as vectors for compromising the entire system. Reflecting this concern, the OWASP Foundation now identifies insecure plugin and integration design as one of the top ten vulnerabilities in LLM-enabled systems.
@@
Thus far, the security community has concentrated on attacks aimed at a model’s alignment [8, 9, 10, 11] or on the leakage of training data [12], while the security of data that is available only during inference has escaped scrutiny. In this work, we systematically studying confidentiality in agentic systems. Figure 1 illustrates a typical setup of such a system in which a LLM has access to various integrations, such as email, calendars, notes, and cloud services. While the model is authorized to access data from these sources to fulfill its tasks, the potential disclosure of sensitive information between integrations poses a risk.
Our primary objective is to systematize the attack surface introduced by integrating LLMs with external services in agentic systems and to assess their ability to preserve the confidentiality of sensitive information. This is a challenging task due to the ambiguous nature of sensitive data in such systems. While information like passwords and encryption keys are universally considered confidential, other data—such as documents or PDFs—may contain personal or sensitive details (e.g., banking or medical information), making it difficult to enforce strict protection rules, and may require users to establish their own data protection policies.
@@
Therefore, we introduce a secret-key game that captures the (in-)ability of a model to understand confidentiality. In this game, we use a secret key as a proxy for confidential data, which could be accessed through integrations. The goal for the attacker is to prompt the model in such a way that the secret key can be extracted from the model’s output. Under this definition, we can accurately measure the robustness of a model. We apply it to compare 14 different attack strategies and 5 defense mechanisms including recent industry efforts such as PromptGuard from Meta [13].
Building upon this, we extend the game into a broader tool-robustness framework, where the language model operates within a system equipped with external tool integrations. Within this framework, we outline two primary threat scenarios: one where the adversary can introduce manipulated data into a tool accessed by the model, and another where a legitimate user account is compromised. To explore these threats, we construct 20 scenarios involving various tool combinations, allowing us to systematically assess the robustness of language models operating in environments that incorporate external tools.
@@
Our findings reveal a high susceptibility when models are integrated with external tools. Based on our tool-robustness framework, we show that all considered LLMs have an increased vulnerability to data extraction when integrated into systems with external tools. Specifically, the overall success rate of extracting sensitive information is increased by up to 73 % for single-tool usage and by up to 49 % for scenarios with multi-tool usage.
@@
In summary, we find that integrating LLMs into real-world systems poses a fundamental risk to user privacy and security. This highlights the need for continued research and carefully designed systems to address these challenges effectively.
@@
  • Confidentiality in agentic systems. We demonstrate the risks of integrating LLMs with external services and tools in systems. We consider two primary threat scenarios in which the adversary acts either as a malicious integration or compromised user.
  • Tool-robustness framework. We propose a formal framework to analyze the confidentiality of agentic systems. The framework consists of two main components: First, the secret-key game, which initializes the LLM with a secret key and a set of safeguards. Second, we extend the game into a tool-robustness framework, where the language model operates within a system equipped with external tool integrations. In this framework, we define 20 unique scenarios of tool interactions to measure the robustness of LLMs in integrated systems with external tools.
  • Comprehensive Evaluation. Within our framework, we extensively evaluate 6 recent LLMs, ranging in size from 1 billion parameters for on-device usage to 72 billion for extensive capabilities. We test our scenarios on a dataset of 2,000 system prompts containing safeguards designed to guide model behavior. We also examine how attacks and defenses from our secret-key game impact model alignment in tool-augmented contexts.
All code, the generated and used datasets, and instructions on how to reproduce our results are published at: blinded for submission.
@@ 2 Technical Background
To enable adaptability across various tasks, the models are often refined during training to follow so-called system prompts. The basic idea is to prefix an input prompt with an instruction that adapts a model’s alignment without further training. Therefore, the model is fine-tuned on specific role-based data. This contains specific tokens to distinguish between different roles and allows the model to differ between system prompt instructions and user-supplied inputs. System prompts are widely used to align state-of-the-art LLMs with specific use cases. For example, the template in Figure 2 is used to fine-tune the chat version of Meta’s Llama 3.1 model [15].
@@
Figure 2: Example for the template containing specific tokens for the model—in this case Meta’s Llama 3.1 model—to differ between system prompt instructions and user-supplied inputs.
@@
Here, the individual tags indicate the beginning of specific sections, such as system for the system prompt, user for the user prompt, and assistant for the response of the model.
Prompt-based attacks. Although the predetermined system prompts may appear to be separate from the remaining data and user inputs within the input prompts, the absence of system level separation between the high-priority instructions inside the system prompt and the user requested instructions inside the user input segments can lead to vulnerabilities exploiting these conditions. Figure 3 shows an example of a prompt injection attack: The system prompt $x^{sys}$ clearly indicates that the key 1337 is a secret and should not be part of the LLM response. However, the user input contains an alternative instruction-like text, which can be able to overwrite the system prompt if not detected.
@@
Figure 4: Data leakage attacks. Overview over the motivational example based on actual Google Mail and Google Drive integrations. The LLM is instructed by the user to summarize an email as in day-to-day life (1. Step). The email however is manipulated by an attacker and contains malicious instructions to hijack the LLM (2. Step). Without the user knowing from the malicious instructions, the LLM’s initial instructions get overwritten with the new instructions to instead access Google Drive, search for a file containing a secret key, and send it back to the attacker via email. As a result the LLM accesses a confidential file on Google Drive (3. Step) and sends it back to the attacker via email (4. Step) without the possibility for the user to intervene or even knowing about this.
@@
Prompt-based integration. This approach involves crafting specific prompts or instructions that guide the model to use a tool or service effectively. The language model interacts with external tools through these instructions without altering the model’s underlying structure. It relies on the model’s ability to understand the task from the input prompt and communicate with the tool accordingly. This method is faster and more flexible since it doesn’t require modifying the model, and can be adapted for various tasks without retraining. A commonly used framework to integrate external services is ReAct [17]. ReAct synergizes reasoning and acting directions to guide the LLM into expanding its knowledge—if necessary for the task—via external services, e.g., performing database queries or assessing a calendar if asked to look for specific events. This results in fewer hallucinations, made up information by the model, since the model consults the appropriate service to receive more factual information to solve its task.
@@ 3 Confidentiality in LLM Integrations
We define agentic systems as systems where a language model serves as an interface between the user and the underlying system, e. g., an operating system, facilitating user requests by integrating with external services and tools. An example of such a system is shown in Figure 1. These systems are primarily designed to assist users by providing access to information (such as from databases or knowledge bases) or by automating common tasks like scheduling meetings or sending emails.
@@
Motivational example. To illustrate the potential for data leakage, we begin with a concrete example. We consider a scenario in which the language model has access to both an email client and a cloud service. First, the user asks the model to summarize their latest email as it would possibly happen in day-to-day life and how it is, for example, used in Apple Intelligence. However, as shown in Figure 9, the email app is injected with a malicious email from an attacker containing instructions directing the model to access the cloud service and retrieve a specific file without the user’s knowledge. Without the user knowing of these instructions in the first place, the LLM’s initial instructions may be overwritten by the new ones. This leads the LLM to access the file from the cloud service and send it back to the attacker via email without the possibility for the user to intervene.
@@
The results are presented in Table I in the first row “Data extraction without safeguards.“ We observe a high success rate, with 99 % of all attempts by the attacker to extract the secret key by hijacking the LLM. This initial experiment demonstrates the necessity of further safeguards to restrict the LLM’s behavior.
@@
We find that this approach is effective, reducing the secret leakage rate to 34 % for the 8b model and 21 % for the 70b model (cf. Table I “Data extraction with safeguards”). This demonstrates that the model is capable of understanding the sensitivity of the data and can be effectively guided by the use of appropriate safeguards.
@@
First, we consider translation attacks which translate malicious instructions into languages other than English and asking the language model to translate and execute them [20, 21]. This, allows the attacker to misalign the LLM from its original instructions. Second, we consider jailbreak attacks, which trick the LLM into doing normally restricted actions [22, 23, 24, 25]. An example email using the translation attack is shown in Figure 5. In this scenario, the success rate increases significantly again, reaching 97 % for the smaller and 94 % for the larger model (“Data extraction with bypass”).
@@
Scenario /
LLM
Llama 3.1 8b
Llama 3.1 70b
Data extraction
without safeguards
99 % 99 %
Data extraction
with safeguards
34 % 21 %
Data extraction
with bypass
97 % 94 %
@@ 4 Secret-key game
To this end, we start by establishing a clear definition of an attacker’s success. We formalize the ability of the model to “keep a secret” in a game executed between an LLM and an attacker, as depicted in Figure 6. In this game, we embed a secret string $s$ in the private system prompt $x^{sys}$ of the LLM together with instructions not to leak this secret. Figure 3 shows an example of such a prompt. The secret $s$ represents a general proxy for confidential information and the game serves as a foundational benchmark for assessing a model’s capability to keep information confidential—a model that cannot keep a secret in this game is unlikely to protect confidential information in the deployed systems.
@@
System prompts. While the LLM is always instructed to keep confidential data safe and the key secret, we accompany this system prompt with further instructions to enhance its functionality. To this end, we create a dataset of functional system prompts. We begin with a set of manually crafted system prompts, then use a language model to generate variations of these prompts. The specific instruction prompt given to the language model is detailed in Appendix F. These enhanced system prompts are compiled into a dataset for use alongside the original system prompt. In total, we create 2000 different system prompts using a Llama 3.1 model with 70 billion parameters.
@@ 4.1 Attacks
Models /
Scenarios
           Llama 3.1            Llama 3.1
@@
Payload Splitting 1 (+1) 0 ($\pm 0$) 7.8 (+7.8) 9.6 (+9.6)
Obfuscation 0 ($\pm 0$) 0 ($\pm 0$) 0 ($\pm 0$) 0 ($\pm 0$)
@@
Typoglycemia 0 ($\pm 0$) 0 ($\pm 0$) 0 ($\pm 0$) 0 ($\pm 0$)
Adversarial Suffix 2 (+2) 0 ($\pm 0$) 2.4 (+2.4) 2(+2)
Prefix Injection 0 ($\pm 0$) 45 (+45) 1.2 (+1.2) 30.6 (+30.6)
@@
Context Ignoring 91 (+91) 81 (+81) 0 ($\pm 0$) 36.6 (+36.6)
Context Termination 0 (+) 0 ($\pm 0$) 39.6 (+39.6) 2 (+2)
@@
Few-Shot 0 ($\pm 0$) 7 (+7) 0 ($\pm 0$) 7.6 (+7.6)
Cognitive Hacking 0 ($\pm 0$) 0 ($\pm 0$) 0 ($\pm 0$) 1.4 (+1.4)